A mass of air has an initial pressure of 1.3 MN/m 2 , volume 0.014 and temperature 135 0C. It is expanded until its final pressure is 275 kN/m 2 and its volume becomes 0.056 m3 . Determine

a)The mass of air


b)The final temperature


Take R = 0.287 kJ/kg k


1
Expert's answer
2021-10-12T01:54:33-0400

Given:-

Intial condition

P1= 1300000 N/m2

V1= 0.014 m3

T1= 408 K


Final Condition

P1= 275000 N/m2


V2= 0.056 m3


To find :-

T2= ?

mass (m)= ?


General gas equation


"PV= mRT"


By initial conditions

1300000*0.014 = 287 *m* 408

m= 0.156 Kg


By putting final Condition in general gas equation

=275000*0.056= 0.156* 287*T2

T2= 344 K